Three years ago Lilly borrowed R10 000 from Faith on condition that she should pay her back two years from now. She also owes Faith R6 000 payable five years from now. The appliable interest rate for both transactions is 13,75% per year,compounded every six months. After considering her payback shedule, Lilly asks Faith if she an pay her R9 000 now and the rest in four years' time. She agrees on ondition that the new agreement will run from now and that an interest rate of 16,28% per year, compounded monthly,

will be appliable from now. The amount that Lilly will have to pay Faith four years from now is


1
Expert's answer
2022-03-14T07:48:14-0400

Lilly owes Faith a total of (10,000+6000)=R16000

If Lilly pays 9000 then she will still owe Faith (16000-9000)= 7000

This amount (7000) is payable in 4 years.

So we want to determine the amount of money that Lilly will pay Faith in four years by compounding the balance of 7000.

Since compounding is done monthly, the interest rate will be divided by 12 while the time will be multiplied by 12 as shown below:

A=P(1+r/12)12t

=7000(1+0.1628/12)4*12

=7000(1.01356)48

=13,362.024

So at the end of four years, Lilly will pay Faith R13,362.024